OpenStudy (anonymous):

Simplify. Remember to use absolute-value notation when necessary. If a root cannot be simplified, state this. √9x²-30x+25

OpenStudy (paxpolaris):

\[\sqrt {\left( a-b \right)^2}= \left| a-b \right|\]

OpenStudy (anonymous):

Yes this how far I got but when I turned it in I got this back

OpenStudy (anonymous):

√9x²-30x+25 9x^2 - 30x + 25 = (3x-5)^2 ; correct Square root of that will be +/- (3x-5). ; incorrect,

OpenStudy (paxpolaris):

\[\sqrt{9x²-30x+25}=\sqrt{\left( 3x \right)^2-2\cdot \left( 3x \right)\cdot5 + 5^2}=\sqrt{\left( 3x-5 \right)^2}\]

OpenStudy (paxpolaris):

the square root symbol only has non-negative roots. eg: \[when\ a^2= 25 \implies a= \pm 5\] \[but\ when\ a=\sqrt{25}; \implies a=+5\]

OpenStudy (paxpolaris):

so, \[\huge\sqrt{\left( 3x-5 \right)^2}= \left| \left( 3x-5 \right) \right|\]

OpenStudy (anonymous):

so that is the finally answer

OpenStudy (paxpolaris):

yea, just with the abs. value sign

OpenStudy (anonymous):

ok thanks that really helped